OpenStudy (anonymous):

what is the limit of lim X--->0 sin^2 * 7t / t

myininaya (myininaya):

\[\lim_{t \rightarrow 0}\frac{\sin^2(7t)}{t}=\lim_{t \rightarrow 0}\frac{\sin(7t) \cdot \sin(7t)}{t}\] \[\lim_{t \rightarrow 0}\frac{\sin(7t)}{t} \cdot \sin(7t)=7 \cdot \lim_{t \rightarrow 0}\frac{\sin(7t)}{7t} \cdot \lim_{t \rightarrow 0}\sin(7t)\]

myininaya (myininaya):

\[=7(1)(\sin(7 \cdot 0))=7\sin(0)=7(0)=0\]

OpenStudy (pokemon23):

i don't understand high level math

OpenStudy (anonymous):

so its 0?

OpenStudy (pokemon23):

all I understand is sin from triangle

OpenStudy (pokemon23):

yes it 0

myininaya (myininaya):

did you really mean x->0?

OpenStudy (pokemon23):

watch the lanaguage

OpenStudy (pokemon23):

read the code of conduct

OpenStudy (anonymous):

lim x→0 x^2 / sin^2 4x

myininaya (myininaya):

yeah just to be careful i wouldn't say that so you also meant sin^2 * 7t or sin^2(7t) because sin^2 * 7t doesn't make sense

OpenStudy (anonymous):

i got it right :) and my bad buddy didnt know sucks was a bad word

myininaya (myininaya):

so this changes thing since x->0 and not t

OpenStudy (pokemon23):

not that...

OpenStudy (pokemon23):

jsut be careful what you say to people..

myininaya (myininaya):

well i think when you put it with that other word its bad

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin^2(7t)}{t}\] so this really was the problem now what I thought it was

OpenStudy (anonymous):

thats one i just did was a new one, that i dont get

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin^2(7t)}{t}=\frac{\sin^2(7t)}{t}\]

myininaya (myininaya):

i thought you meant t->0 earlier

myininaya (myininaya):

now that i know you actually meant x->0 this is correct

OpenStudy (anonymous):

the one i did earlier had 7t sin^2.... but i just posted one that was like lim x→0 x^2 / sin^2 4x

myininaya (myininaya):

i'm confused

myininaya (myininaya):

which one did you mean? x->0 or t->0

OpenStudy (anonymous):

i need help with the last one where3 x -->0

myininaya (myininaya):

i been talking about the first one you put up

myininaya (myininaya):

t->0 or x->0?

OpenStudy (anonymous):

the first one is right you did great!

myininaya (myininaya):

you said x->0 i did the problem for t->0 not x->0

myininaya (myininaya):